# ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection

### v1

@@ Abstract
Zero-shot anomaly detection (ZSAD) aims to detect anomalies without any target domain training samples, relying solely on external auxiliary data. Existing CLIP-based methods attempt to activate the model’s ZSAD potential via handcrafted or static learnable prompts. The former incur high engineering costs and limited semantic coverage, whereas the latter apply identical descriptions across diverse anomaly types, thus fail to adapt to complex variations. Furthermore, since CLIP is originally pretrained on large-scale classification tasks, its anomaly segmentation quality is highly sensitive to the exact wording of class names, severely constraining prompting strategies that depend on class labels. To address these challenges, we introduce ViP2-CLIP. The key insight of ViP2-CLIP is a Visual-Perception Prompting (ViP-Prompt) mechanism, which fuses global and multi-scale local visual context to adaptively generate fine-grained textual prompts, eliminating manual templates and class-name priors. This design enables our model to focus on precise abnormal regions, making it particularly valuable when category labels are ambiguous or privacy-constrained. Extensive experiments on 15 industrial and medical benchmarks demonstrate that ViP2-CLIP achieves state-of-the-art performance and robust cross-domain generalization.
ViP2-CLIP: Visual-Perception Prompting with Unified Alignment for Zero-Shot Anomaly Detection
Ziteng Yang1,*    JingzehuaXu1,*    Yanshu Li2    Zepeng Li1 Yeqiang Wang3    Xinghui Li1$\dagger$ 1Tsinghua Shenzhen International Graduate School 2Brown University    3Northwest A&F University
@@ 1 Introduction
Large-scale vision–language models (VLMs) such as CLIP (Radford et al., 2021) have shown impressive zero-shot recognition capabilities by pre-training on vast image–text pairs. Inspired by this success, CLIP has been widely adopted for Zero-Shot Anomaly Detection (ZSAD) (Jeong et al., 2023), which seeks to detect anomalies automatically without any target domain training samples, offering a practical solution for data-scarce scenarios. Unlike conventional image classification, which focuses on foreground semantics, anomaly detection emphasizes irregular regions on object surfaces. Consequently, existing CLIP-based ZSAD methods typically construct two sets of text prompts—‘normal’ and ‘anomalous’, and align them with image features in a shared semantic space. As illustrated in Figure 1(a), text prompts are aligned with the global image feature to infer image-level anomaly score, while alignment with local patch features then generates pixel-level anomaly maps.
@@
Currently, state-of-the-art ZSAD methods focus on prompt engineering to leverage CLIP’s generalization. Representative approaches such as WinCLIP (Jeong et al., 2023), AnomalyCLIP (Zhou et al., 2024), and VCP-CLIP (Qu et al., 2024) either craft semantically rich manual templates or learn prompt tokens to boost performance. However, manual templates demand extensive human effort and cannot guarantee comprehensive semantic coverage. Although some works (Zhou et al., 2024) introduce object-agnostic learnable prompts to capture generic normal and abnormal concepts, they apply the same description across diverse anomaly types, hindering adaptability to complex semantic variations. Moreover, methods like CLIP-AD (Chen et al., 2024), AdaCLIP (Cao et al., 2024), FILo (Gu et al., 2024), and KANOCLIP (Li et al., 2025) rely heavily on class names as prompt priors. Yet we find that CLIP’s segmentation quality remains acutely sensitive to class-name variations, even after fine-tuning on auxiliary anomaly data (see Appendix A for detailed analysis). This likely stems from the limited scope of downstream adaptation, which cannot fully stabilize local alignment against label shifts. Such sensitivity constrains the class-name usage in prompts, highlighting the urgent need for robust, class-name agnostic but adaptive prompting strategies in ZSAD.
@@
To this end, we propose a Visual-Perception Prompting (ViP-Prompt) mechanism composed of an Image-Conditioned Adapter (ICA) and a Fine-Grained Perception Module (FGP). ICA adaptively injects global visual context into the prompts’ embedding space, while FGP integrates multi-scale patch features, enabling the prompts to capture fine-grained irregularities. By replacing fixed class-name tokens with image-conditioned prompts that fuse global and local cues, ViP-Prompt significantly enhances prompts’ robustness and reinforces cross-modal alignment across diverse semantic scenarios. Building upon that, we further introduce ViP2-CLIP, a ZSAD model that integrates visual-perception prompts with a unified alignment mechanism; its overall architecture is depicted in Figure 1(b). ViP2-CLIP first employs ViP-Prompt to condition text prompts on both global and local visual cues, fully leveraging CLIP’s ZSAD performance; it then adopts a Unified Text-Patch Alignment (UTPA) strategy to align prompt tokens with multi-scale patch features through a single optimization, yielding superior precision in both detection and localization.
Overall, our key contributions are summarized as follows:
@@
  • We propose ViP-Prompt, which fuses global and local visual features to generate image-conditioned normal and abnormal prompts, fully exploiting CLIP’s potential for ZSAD. This design eliminates the need for handcrafted templates or class-name priors, delivers stronger generalization and robustness, making it particularly valuable when category labels are ambiguous or privacy-constrained.
@@
  • Comprehensive experiments on 15 real-world industrial and medical datasets demonstrate that ViP2-CLIP consistently achieves superior ZSAD performance.
@@ 2 Related Work
Zero-shot anomaly detection (ZSAD) localizes abnormal regions without any target class samples during training. The advent of CLIP (Radford et al., 2021) has spurred notable advances in this task. WinCLIP (Jeong et al., 2023) first ported CLIP to ZSAD with diverse handcrafted prompts. However, since CLIP is pretrained to align with general object semantics rather than abnormal patterns, its anomaly detection capability remains limited. To ease this burden, APRIL-GAN (Chen et al., 2023) and CLIP-AD (Chen et al., 2024) append shallow linear heads and fine-tune them on auxiliary datasets. AnomalyCLIP (Zhou et al., 2024) removes handcrafted templates via learnable, object-agnostic prompts, boosting cross-domain generalization, while AdaCLIP (Cao et al., 2024) employs hybrid prompts to align text and image features better. VCP-CLIP (Qu et al., 2024) instead projects the image’s global feature into a fixed handcrafted prompt space, which enhances segmentation performance. By contrast, our ViP-Prompt fuses multi-scale visual cues with a learnable seed to generate dynamic, manual-free prompts, fully exploiting CLIP’s ZSAD performance, boosting both effectiveness and robustness across diverse domains.
@@
Prompt learning was initially proposed in natural language processing to enhance the adaptability of pre-trained models (Ouyang et al., 2022; Touvron et al., 2023) to diverse downstream tasks. CoOp (Radford et al., 2021) first brought this paradigm to the vision domain by inserting trainable tokens into text inputs, allowing CLIP to adapt to specific tasks without fine-tuning. Static prompts, however, soon proved brittle on unseen classes. To enhance transferability, CoCoOp (Zhou et al., 2022) and DenseCLIP (Rao et al., 2022) generate image-conditioned prompts that respond to visual context. Later works such as MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) push prompting into both modalities, jointly tuning image and text space for more substantial cross-modal alignment. More recent works (Yao et al., 2023, 2024) enforce prompt generalization by regularizing learnable tokens toward handcrafted templates, reducing forgetting and improving unseen class recognition.
@@ 3 Preliminary
Unlike conventional classification, ZSAD flags deviations from normality rather than assigning foreground semantics. Most approaches, therefore, instantiate two text prompts Jeong et al. (2023), a normal prompt $g_{n}$ and an abnormal prompt $g_{a}$. At the image level, they compare the global image feature $f_{i}$ with both prompts and take $P(g_{a},f_{i})$ as the anomaly score; while at the pixel level, for each location $(j,k)$, they extract the patch token $f_{i}^{m}(j,k)$, and compute the local normal score $S_{n}(j,k)=P\bigl{(}g_{n},f_{i}^{m}(j,k)\bigr{)}$ and anomaly score $S_{a}(j,k)=P\bigl{(}g_{a},f_{i}^{m}(j,k)\bigr{)}$, from which a pixel-wise anomaly map is generated.
@@ 4 Method
We propose ViP2-CLIP, which leverages Visual-Perception Prompting (ViP-Prompt) to fully exploit CLIP’s ZSAD capabilities. As illustrated in Figure 2, ViP2-CLIP first introduces ViP-Prompt to fuse learnable normal and abnormal prompts with both global and multi-scale local visual features. This design enables the prompts to adaptively capture the fine-grained visual patterns of the inspected image, thereby improving cross-modal semantic alignment. Furthermore, we adopt a Unified Text-Patch Alignment (UTPA) scoring strategy, which optimizes alignment between prompt tokens and multi-scale patch features to support anomaly detection and localization jointly. By aggregating alignment signals across multiple layers, ViP2-CLIP produces robust anomaly scores and fine-grained anomaly maps.
@@ 4.1 Visual-Perception Prompt (ViP-Prompt)
Effective prompt design plays a crucial role in achieving strong ZSAD performance. Handcrafted templates (Jeong et al., 2023; Chen et al., 2023) are costly and fragile, while static learnable prompts (Zhou et al., 2024) apply identical textual descriptions to all detection targets, limiting their ability to adapt to complex semantic variations. Furthermore, most existing techniques (Chen et al., 2024; Cao et al., 2024; Gu et al., 2024; Li et al., 2025) generate prompts based on class names, since CLIP’s anomaly segmentation quality is highly sensitive to class-name wording, these methods remain unusable in label-ambiguous scenarios. To overcome these limitations, we therefore introduce ViP-Prompt, which is built from an Image-Conditioned Adapter (ICA) and Fine-Grained Perception Module (FGP). By fusing learnable tokens with the global and local visual context, ViP-Prompt generates multi-level descriptions that flexibly track objects’ visual patterns, thus markedly exploiting CLIP’s ZSAD performance in diverse scenarios.
@@
To condition prompts on the target object and thus adapt to complex detect scenarios, we map the global visual embedding $f$ extracted by the image encoder through a lightweight Meta-Net $h_{i}(\cdot)$ (Linear–ReLU–Linear) into the text embedding space, denoted as $\{a_{i}\}_{i=1}^{r}=h_{i}(f)$, where $r$ is the number of mapped tokens and $a_{i}\in\mathbb{R}^{C}$, then fuse it with static learnable vector to generate dynamic prompt token. Thus, we define the prompts’ structure as follows:
@@
where $z_{i}=o_{i}+a_{i}$. By projecting the image’s global feature into text embedding space, ICA adaptively generates normal and anomalous descriptors conditioned on the target object, eliminating the dependence on explicit class labels and enabling more robust cross-modal alignment in privacy-constrained settings.
@@
To assess FGP’s fine-grained sensitivity, Figure 3 visualizes the attention maps of both normal and abnormal prompts across diverse images. Normal prompts activate on regular textures and structures of the image, while anomalous prompts concentrate on defects and irregular regions, indicating high sensitivity to anomalous patterns. These results demonstrate that FGP effectively refines prompt embeddings by incorporating localized visual cues, thereby enhancing cross-modal alignment.
@@ 4.2 Unified Text-Patch Alignment (UTPA)
Recent CLIP-based ZSAD models typically adopt a dual-branch alignment scheme Chen et al. (2024); Zhou et al. (2024); Chen et al. (2023): one branch aligns the text prompt with global visual embedding to score anomalies at the image level, while the other aligns it with local patch embeddings for pixel-level anomaly maps. However, because these branches operate in separate feature subspaces, it hinders the simultaneous attainment of high-precision anomaly detection and localization.
To address this limitation, we introduce the UTPA strategy, which treats anomaly detection and localization as a unified alignment task between text embeddings and local patch embeddings. Concretely, to obtain the pixel-level anomaly map, we follow the standard practice: for the $l$-th layer feature map $F_{m}^{l}$, the patch token at spatial location $(j,k)$, denoted as $f_{j,k}^{l}\in\mathbb{R}^{D}$, receives its anomaly score by computing the similarity $P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}$ as follows:
${}{P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}=\frac{\exp\!\bigl{(}\langle\tilde{g_{a}},\,f_{j,k}^{l}\rangle/\tau\bigr{)}}{\exp\!\bigl{(}\langle\tilde{g_{n}},\,f_{j,k}^{l}\rangle/\tau\bigr{)}+\exp\!\bigl{(}\langle\tilde{g_{a}},\,f_{j,k}^{l}\rangle/\tau\bigr{)}}}.$ (7)
@@
where $Up(\cdot)$ denotes channel-wise upsampling.
Unlike prior works (Chen et al., 2024; Zhou et al., 2024; Chen et al., 2023) that rely on global alignment, we contend that an image’s abnormality is driven by its most defective regions. Consequently, our image-level score is still derived from local patch features. For the $l$-th layer feature map $F_{m}^{l}$, we select the Top-k patch tokens with the highest $P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}$ among all patches. The image’s representative anomaly feature $F_{a}^{l}$ is then obtained via global average pooling over this subset:
@@
where $\mathcal{T}_{l}$ indexes the chosen Top-k patches. The Top-k pooling directs the model’s attention to the most suspicious patches, ensuring that the subsequent anomaly score faithfully reflects the severity of localized defects. Consequently, the image-level anomaly score $S_{a}^{l}$ at the $l$-th layer is obtained by computing $S_{a}^{l}=P\bigl{(}\tilde{g_{a}},\,F_{a}^{l}\bigr{)}$.
4.3 Training and Inference
During training, ViP2-CLIP is optimized by minimizing a global loss $L_{\mathrm{global}}$ and a local loss $L_{\mathrm{local}}$:
here, $\lambda$ is a hyperparameter to balance the global and local losses; $N$ represents the number of intermediate layers. $L_{\mathrm{global}}$ is the cross-entropy loss matching the cosine similarity between text embeddings $\tilde{g_{n}}/\tilde{g_{a}}$ and the representative anomalous visual feature $F_{a}^{l}$; $L_{\mathrm{local}}$ is computed as the sum of a Focal loss Lin et al. (2017) and a Dice loss Li et al. (2019) to optimize local alignment jointly:
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes channel-wise concatenation; for each layer $l$, $M_{n}^{l}$ and $M_{a}^{l}$ are the normal and anomaly score maps; $S_{\mathrm{gt}}$ is the ground-truth mask; and $I$ is the all-ones matrix.
During inference, the final image-level anomaly score is computed as:$\mathrm{Score}=\frac{1}{N}\sum_{l=1}^{N}S_{a}^{l}$. For pixel-level prediction, we merge all intermediate maps $M_{n}^{l}$ and $M_{a}^{l}$, then obtain the anomaly map $\mathrm{Map}\in\mathbb{R}^{H\times W}$ as follows:
${\mathrm{Map}=G_{\sigma}\!\Biggl{(}\frac{1}{N}\sum_{l=1}^{N}\Bigl{(}\tfrac{1}{2}\bigl{(}I-\mathrm{Up}(M_{n}^{l})\bigr{)}+\tfrac{1}{2}\,\mathrm{Up}(M_{a}^{l})\Bigr{)}\Biggr{)}},$ (12)
where $G_{\sigma}$ denotes Gaussian smoothing.
5 Experiments
5.1 Setup
Datasets & Baselines
To comprehensively evaluate ViP2-CLIP across diverse application scenarios, we conduct extensive experiments on public benchmarks spanning industrial and medical domains. In the industrial domain, we use MVTec AD (Bergmann et al., 2019), VisA (Zou et al., 2022), MPDD (Jezek et al., 2021), BTAD (Mishra et al., 2021), KSDD (Tabernik et al., 2020), DAGM (Wieler and Hahn, 2007), and DTD-Synthetic (Aota et al., 2023). For medical imaging, we include brain tumour detection datasets HeadCT (Salehi et al., 2021), BrainMRI (Salehi et al., 2021), and Br35H (Hamada, 2020), skin cancer detection dataset ISIC Codella et al. (2018), colon polyp detection datasets CVC-ClinicDB Bernal et al. (2015), CVC-ColonDB (Tajbakhsh et al., 2015), Endo (Hicks et al., 2021), and Kvasir (Jha et al., 2019). We compare ViP2-CLIP against five leading ZSAD methods: CLIP (Radford et al., 2021), WinCLIP (Jeong et al., 2023), APRIL-GAN Chen et al. (2023), AnomalyCLIP (Zhou et al., 2024) and AdaCLIP (Cao et al., 2024). More details about the datasets or methods are shown in Appendix B.
@@ 5.1 Setup
Table 1: ZSAD performance comparison on industrial domain. * denotes results taken from original papers. The best performance is highlighted in red, and the second in blue.
@@
Table 2: ZSAD performance comparison on medical domain. The best performance is highlighted in red, and the second in blue. Note that the image-level medical AD datasets do not provide pixel-level segmentation annotations, so the pixel-level medical AD datasets are different from the image-level datasets.

### v2

@@ Abstract
Zero‑shot anomaly detection (ZSAD) aims to detect anomalies without any target domain training samples, relying solely on external auxiliary data. Existing CLIP-based methods attempt to activate the model’s ZSAD potential via handcrafted or static learnable prompts. The former incur high engineering costs and limited semantic coverage, whereas the latter apply identical descriptions across diverse anomaly types, thus fail to adapt to complex variations. Furthermore, since CLIP is originally pretrained on large-scale classification tasks, its anomaly segmentation quality is highly sensitive to the exact wording of class names, severely constraining prompting strategies that depend on class labels. To address these challenges, we introduce ViP2-CLIP. The key insight of ViP2-CLIP is a Visual‑Perception Prompting (ViP‑Prompt) mechanism, which fuses global and multi‑scale local visual context to adaptively generate fine‑grained textual prompts, eliminating manual templates and class-name priors. This design enables our model to focus on precise abnormal regions, making it particularly valuable when category labels are ambiguous or privacy-constrained. Extensive experiments on 15 industrial and medical benchmarks demonstrate that ViP2-CLIP achieves state‑of‑the‑art performance and robust cross‑domain generalization.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
@@
Large‑scale vision–language models (VLMs) such as CLIP [23] have shown impressive zero‑shot recognition capabilities by pre‑training on vast image–text pairs. Inspired by this success, CLIP has been widely adopted for Zero‐Shot Anomaly Detection (ZSAD) [12], which seeks to detect anomalies automatically without any target domain training samples, offering a practical solution for data‑scarce scenarios. Unlike conventional image classification, which focuses on foreground semantics, anomaly detection emphasizes irregular regions on object surfaces. Consequently, existing CLIP‑based ZSAD methods typically construct two sets of text prompts—‘normal’ and ‘anomalous’, and align them with image features in a shared semantic space. As illustrated in Fig. 1 (a), text prompts are aligned with the global image feature to infer image-level anomaly score, while alignment with local patch features then generates pixel-level anomaly maps.
@@
Currently, state-of-the-art ZSAD methods focus on prompt engineering to leverage CLIP’s generalization. Representative approaches such as WinCLIP [12], AnomalyCLIP [33], and VCP-CLIP [22] either craft semantically rich manual templates or learn prompt tokens to boost performance. However, manual templates demand extensive human effort and cannot guarantee comprehensive semantic coverage. Although some works [33] introduce object-agnostic learnable prompts to capture generic normal and abnormal concepts, they apply the same description across diverse anomaly types, hindering adaptability to complex semantic variations. Moreover, methods like CLIP-AD [6], AdaCLIP [4], FILo [8], and KANOCLIP [17] rely heavily on class names as prompt priors. Yet we find that CLIP’s segmentation quality remains acutely sensitive to semantically equivalent class-name variations, even after fine-tuning on auxiliary anomaly data (see Appendix A for detailed analysis). This likely stems from the limited scope of downstream adaptation, which cannot fully stabilize local alignment against label shifts. Such sensitivity constrains the class-name usage in prompts, highlighting the urgent need for robust, class-name agnostic but adaptive prompting strategies in ZSAD.
@@
To this end, we propose a Visual‑Perception Prompting (ViP‑Prompt) mechanism composed of an Image‑Conditioned Adapter (ICA) and a Fine‑Grained Perception Module (FGP). ICA adaptively injects global visual context into the prompts’ embedding space, while FGP integrates multi‑scale patch features, enabling the prompts to capture fine‑grained irregularities. By replacing fixed class-name tokens with image-conditioned prompts that fuse global and local cues, ViP-Prompt significantly enhances prompts’ robustness and reinforces cross-modal alignment across diverse semantic scenarios. Building upon that, we further introduce ViP2-CLIP, a ZSAD model that integrates visual-perception prompts with a unified alignment mechanism; its overall architecture is depicted in Fig. 1 (b). ViP2-CLIP first employs ViP-Prompt to condition text prompts on both global and local visual cues, fully leveraging CLIP’s ZSAD performance; it then adopts a Unified Text-Patch Alignment (UTPA) strategy to align prompt tokens with multi-scale patch features through a single optimization, yielding superior precision in both detection and localization.
Our key contributions are summarized as follows:
@@
  • We propose ViP-Prompt, which fuses global and local visual features to generate image-conditioned prompts, fully exploiting CLIP’s potential for ZSAD. This design eliminates the need for handcrafted templates or class-name priors, delivers stronger generalization and robustness, making it particularly valuable when category labels are ambiguous or privacy‑constrained.
@@
  • Comprehensive experiments on 15 industrial and medical datasets demonstrate that ViP2-CLIP consistently achieves superior ZSAD performance.
@@ 2 Related Work
ZSAD localizes abnormal regions without any target class samples during training. The advent of CLIP [23] has spurred notable advances in this task. WinCLIP [12] first ported CLIP to ZSAD with diverse handcrafted prompts. However, since CLIP is pretrained to align with general object semantics rather than abnormal patterns, its anomaly detection capability remains limited. To ease this burden, APRIL-GAN [5] and CLIP-AD [6] append shallow linear heads and fine-tune them on auxiliary datasets. AnomalyCLIP [33] removes handcrafted templates via learnable, object-agnostic prompts, boosting cross-domain generalization, while AdaCLIP [4] employs hybrid prompts to align text and image features better. VCP-CLIP [22] instead projects the image’s global feature into a fixed handcrafted prompt space, which enhances segmentation performance. By contrast, our ViP-Prompt fuses multi-scale visual cues with a learnable seed to generate dynamic, manual-free prompts, fully exploiting CLIP’s ZSAD performance, boosting both effectiveness and robustness in anomaly detection and localization.
@@
Prompt learning was initially proposed in natural language processing to enhance the adaptability of pre-trained models [21, 28] to diverse downstream tasks. CoOp [23] first brought this paradigm to the vision domain by inserting trainable tokens into text inputs, allowing CLIP to adapt to specific tasks without fine‑tuning. Static prompts, however, soon proved brittle on unseen classes. To enhance transferability, CoCoOp [32] and DenseCLIP [24] generate image-conditioned prompts that respond to visual context. Recent works [30, 31] enforce prompt generalization by regularizing learnable tokens toward handcrafted templates, improving unseen class recognition. Works as MaPL [15], PromptSRC [16] and MMRL [9] push prompting into both modalities, jointly tuning image and text space for more substantial cross-modal alignment.
@@ 3 Preliminary
Unlike conventional classification, ZSAD flags deviations from normality rather than assigning foreground semantics. Most approaches, therefore, instantiate two text prompts [12], a normal prompt $g_{n}$ and an abnormal prompt $g_{a}$. At the image level, they compare the global image feature $f_{i}$ with both prompts and take $P(g_{a},f_{i})$ as the anomaly score; while at the pixel level, for each location $(j,k)$, they extract the patch token $f_{i}^{m}(j,k)$, and compute the local normal score $S_{n}(j,k)=P\bigl{(}g_{n},f_{i}^{m}(j,k)\bigr{)}$ and anomaly score $S_{a}(j,k)=P\bigl{(}g_{a},f_{i}^{m}(j,k)\bigr{)}$, from which a pixel-wise anomaly map is generated.
@@ 4 Method
We propose ViP2-CLIP, which leverages Visual-Perception Prompting (ViP-Prompt) to fully exploit CLIP’s ZSAD capabilities. As illustrated in Fig. 2, ViP2-CLIP first introduces ViP-Prompt (Sec. 4.1) to fuse learnable normal and abnormal prompts with both global and multi-scale local visual features. This design enables prompts to adaptively capture the fine‑grained visual patterns of the inspected image, thereby improving cross-modal semantic alignment. Furthermore, we adopt a Unified Text-Patch Alignment (UTPA) scoring strategy (Sec. 4.2), which optimizes alignment between prompt tokens and multi-scale patch features to support anomaly detection and localization jointly. By aggregating alignment signals across multiple layers, ViP2-CLIP produces robust anomaly scores and fine-grained anomaly maps.
@@ 4.1 Visual-Perception Prompt (ViP-Prompt)
We propose ViP-Prompt, an adaptive fine-grained prompting framework, which is built from an Image-Conditioned Adapter (ICA) and Fine-Grained Perception Module (FGP). By fusing learnable tokens with the global and local visual context, ViP-Prompt generates multi-level descriptions that flexibly track objects’ visual patterns, thus markedly exploiting CLIP’s ZSAD performance in diverse scenarios.
@@
To condition prompts on the target object and thus adapt to complex detectionl scenarios, we map the global visual embedding $f$ extracted by the image encoder through a lightweight Meta-Net $h_{i}(\cdot)$ (Linear–ReLU–Linear) into the text embedding space, denoted as $\{a_{i}\}_{i=1}^{r}=h_{i}(f)$, where $r$ is the number of mapped tokens and $a_{i}\in\mathbb{R}^{C}$, then fuse it with static learnable vector to generate dynamic prompt token. Thus, we define the prompts’ structure as follows:
@@
where $z_{i}=o_{i}+a_{i}$. By projecting the image’s global feature into text embedding space, ICA adaptively generates normal and anomalous descriptors conditioned on the target object, eliminating the dependence on explicit class labels and enabling more robust cross‑modal alignment in privacy‑constrained settings.
@@
To assess FGP’s fine-grained sensitivity, Fig. 3 visualizes the attention maps of both normal and abnormal prompts across diverse images. Normal prompts activate on regular textures and structures of the image, while anomalous prompts concentrate on defects and irregular regions, indicating high sensitivity to anomalous patterns. These results demonstrate that FGP effectively refines prompt embeddings by incorporating localized visual cues, thereby enhancing cross-modal alignment.
@@ 4.2 Unified Text-Patch Alignment (UTPA)
Recent CLIP-based ZSAD models typically adopt a dual-branch alignment scheme [6, 33, 5]: one branch aligns the text prompt with global visual embedding to score anomalies at the image level, while the other aligns it with local patch embeddings for pixel-level anomaly maps. However, because these branches operate in separate feature subspaces, it hinders the simultaneous attainment of high-precision anomaly detection and localization.
To address this limitation, we introduce the UTPA strategy, which treats anomaly detection and localization as a unified alignment task between text embeddings and local patch embeddings. Concretely, to obtain the pixel-level anomaly map, we follow the standard practice: for the $l$-th layer feature map $F_{m}^{l}$, the patch token at spatial location $(j,k)$, denoted as $f_{j,k}^{l}\in\mathbb{R}^{D}$, receives its anomaly score by computing the similarity $P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}$, as defined in Eq. 1.
@@
where $Up(\cdot)$ denotes channel‐wise upsampling.
Unlike prior works [6, 33, 5] that rely on global alignment, we contend that an image’s abnormality is driven by its most defective regions. Consequently, our image-level score is still derived from local patch features. For the $l$-th layer feature map $F_{m}^{l}$, we select the Top-k patch tokens with the highest $P\bigl{(}\tilde{g_{a}},\,f_{j,k}^{l}\bigr{)}$ among all patches. The image’s representative anomaly feature $F_{a}^{l}$ is then obtained via global average pooling over this subset:
@@
where $\mathcal{T}_{l}$ indexes the chosen Top-k patches. The Top-k pooling directs the model’s attention to the most suspicious patches, ensuring that the subsequent anomaly score faithfully reflects the severity of localized defects. Consequently, the image-level anomaly score $S_{a}^{l}$ at the $l$-th layer is obtained by computing $S_{a}^{l}=P\bigl{(}\tilde{g_{a}},\,F_{a}^{l}\bigr{)}$. By enforcing a unified text-patch alignment, UTPA effectively overcomes the optimization conflict of traditional dual-branch schemes, yielding more balanced ZSAD performance, as further validated in Sec. 5.3.
@@
Table 1: ZSAD performance comparison on industrial domain. * denotes results taken from original papers. The best performance is shown in red, with the second-best highlighted in blue.
@@
Table 2: ZSAD performance comparison on medical domain. The best performance is shown in red, with the second-best highlighted in blue. Note that the image-level medical AD datasets do not provide pixel-level segmentation annotations, so the pixel-level medical AD datasets are different from the image-level datasets.
4.3 Training and Inference
During training, ViP2-CLIP is optimized by minimizing a global loss $L_{\mathrm{global}}$ and a local loss $L_{\mathrm{local}}$ [33]:
here, $\lambda$ is a hyperparameter to balance the global and local losses; $N$ represents the number of intermediate layers. $L_{\mathrm{global}}$ is the cross‐entropy loss matching the cosine similarity between text embeddings $\tilde{g_{n}}/\tilde{g_{a}}$ and the representative anomalous visual feature $F_{a}^{l}$; $L_{\mathrm{local}}$ is computed as the sum of a Focal loss [19] and a Dice loss [18] to optimize local alignment jointly:
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes channel‐wise concatenation; for each layer $l$, $M_{n}^{l}$ and $M_{a}^{l}$ are the normal and anomaly score maps; $S_{\mathrm{gt}}$ is the ground‐truth mask; and $I$ is the all‐ones matrix.
During inference, the final image‐level anomaly score is computed as:$\mathrm{Score}=\frac{1}{N}\sum_{l=1}^{N}S_{a}^{l}$. For pixel‐level prediction, we merge all intermediate maps $M_{n}^{l}$ and $M_{a}^{l}$, then obtain the anomaly map $\mathrm{Map}\in\mathbb{R}^{H\times W}$ as follows:
$\mathrm{Map}=G_{\sigma}\!\Biggl{(}\frac{1}{N}\sum_{l=1}^{N}\Bigl{(}\tfrac{1}{2}\bigl{(}I-\mathrm{Up}(M_{n}^{l})\bigr{)}+\tfrac{1}{2}\,\mathrm{Up}(M_{a}^{l})\Bigr{)}\Biggr{)}$ (11)
where $G_{\sigma}$ denotes Gaussian smoothing.
5 Experiments
5.1 Setup
Datasets & Baselines
To comprehensively evaluate ViP2-CLIP across diverse application scenarios, we conduct extensive experiments on public benchmarks spanning industrial and medical domains. In the industrial domain, we use MVTec AD [2], VisA [34], MPDD [13], BTAD [20], KSDD [26], DAGM [29], and DTD-Synthetic [1]. For medical imaging, we include brain tumour detection datasets HeadCT [25], BrainMRI [25], and Br35H [10], skin cancer detection dataset ISIC [7], colon polyp detection datasets CVC-ClinicDB [3], CVC-ColonDB [27], Endo [11], and Kvasir [14]. We compare ViP2-CLIP against five leading ZSAD methods: CLIP [23], WinCLIP [12], APRIL-GAN [5], AnomalyCLIP [33] and AdaCLIP [4]. More details about the datasets or methods are shown in Appendix B.